The length of a rectangle is 10 more than 3 times its width. The perimeter of the rectangle is 108. Find the dimensions of the rectangles.
Answer by Paul(988)

Let x be the width
2[x+(3x+10)]=108
2(4x+10)=108
8x+20=108
8x=88
x=11
3(11)+10=43
Hence, the width is 11 and the length is 43